OpenStudy (anonymous):

A 1000L brine tank is filled with 15kg of salt. The tank is constantly mixed. Water enters and exits at 10L/min. Set up a differential equation for the amount of salt at any time.

OpenStudy (anonymous):

Give me a minute

OpenStudy (anonymous):

dy/dt = (rate in) - (rate out)... rate in = 0 (only water is entering the tank) rate out = (y(t)/1000) kg/L * 10 L/min = y(t)/100 kg/min... dy/dt = -y/100

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

Shouldn't t be in the equation?

OpenStudy (anonymous):

dy/dt = -y(t)/100

OpenStudy (anonymous):

It is correct how it is now... What you'll have to do now is just integrate by separation of variables or by using the integrating factor.

OpenStudy (anonymous):

Using an integrating factor may be the easiest route. Don't forget to use the initial condition: y(0) = 15kg to obtain the particular solution.

OpenStudy (anonymous):

So if \[y \prime(t) = -y(t)/100\] then \[y(t)=-1/100t + C\]?

OpenStudy (anonymous):

No... y(t) = ce^(-t/100) So when t = 10, The particular solution will be: y = 15e^(-t/100)

OpenStudy (anonymous):

Sorry, when t = 0

OpenStudy (anonymous):

Ah, okay. Gotcha. Thank you very much.